a. Find the value of z subzero such that P(z > zsubzero) = .5

b.  Find the value of z subzero such that P(z < zsubzero) = .8643

c. Find the value of z subzero such that P(-z subzero < Z < zsubzero) = .9

d. Find the value of z subzero such that P(-z subzero < Z < zsubzero) = .99
Answer #1

Solution :

a) P(Z > z ) = 0.50

1- P(z < z) =0.50

P(z < z) = 1-0.50 = 0.50

z = 0

b)

P(Z < z) = 0.8643

z = 1.1

c)

P(-z \leq Z \leq z) = 0.90

P(Z \leq z) - P(Z \leq -z) = 0.90

2P(Z \leq z) - 1 = 0.90

2P(Z \leq z) = 1 + 0.90 = 1.90

P(Z \leq z) = 1.90 / 2 = 0.95

P(Z \leq \pm 1.645) = 0.95

z = - 1.645 , +1.645

d)

P(-z \leq Z \leq z) = 0.99

P(Z \leq z) - P(Z \leq -z) = 0.99

2P(Z \leq z) - 1 = 0.99

2P(Z \leq z) = 1 + 0.99 = 1.995

P(Z \leq z) = 1.99 / 2 = 0.995

P(Z \leq \pm 2.576) = 0.995

z = - 2.576 , +2.576